OpenStudy (anonymous):

write the equation (in standard form) of the line with m=4 and b=6

jimthompson5910 (jim_thompson5910):

y = mx+b y = 4x+6 y - 4x = 6 -4x + y = 6 4x - y = -6

OpenStudy (anonymous):

isnt standard form Ax+By+C=0 though?

jimthompson5910 (jim_thompson5910):

if you want it in that form then add 6 to both sides to get 4x - y + 6 = 0 standard form has 2 versions

OpenStudy (anonymous):

in my answers it says 0=4x-y+6 so is that Ax+By+C=0 version?

OpenStudy (anonymous):

in my answers it says 0=4x-y+6 so is that Ax+By+C=0 version?

jimthompson5910 (jim_thompson5910):

close enough, it's just flipped around

OpenStudy (anonymous):

how did u add 6 to both sides?

jimthompson5910 (jim_thompson5910):

I added 6 to the right side to get -6+6 = 0 and 6 to the left side to balance it out